# Chlorophyll-a Mapping and Prediction in the Mar Menor Lagoon Using C2RCC-Processed Sentinel 2 Imagery

### v1

@@ V-A Results
When comparing the performance obtained in the upper layers with the studies summarized in Table II our metrics are on par with previous findings for the Mar Menor. However, it is worth noting that our dataset spans nearly a decade, covers the widest range of Chl-a concentrations reported so far, offers the highest spatial resolution, and includes a complete validation framework. These factors make the results particularly robust, albeit at the expense of a more complex and computationally demanding processing workflow.
@@
TABLE VI: Test $R^{2}$ for depth 1-2 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
C2X-Complex_rhow_5x5_depth_in_1_2 0.86 0.60 0.87 0.77 0.82 0.71 0.76 0.73 0.86 0.83
C2X-Complex_rhow_9x9_depth_in_1_2 0.86 0.55 0.84 0.76 0.80 0.53 0.66 0.79 0.83 0.83
C2X-Complex_rhown_5x5_depth_in_1_2 0.86 0.63 0.86 0.78 0.80 0.71 0.74 0.71 0.86 0.78
C2X_rhow_3x3_depth_in_1_2 0.82 0.71 0.62 0.75 0.84 0.67 0.75 0.77 0.63 0.83
C2X-Complex_rhow_3x3_depth_in_1_2 0.80 0.56 0.80 0.79 0.76 0.54 0.72 0.69 0.83 0.77
C2X-Complex_rhown_3x3_depth_in_1_2 0.82 0.59 0.83 0.75 0.74 0.67 0.74 0.70 0.78 0.77
C2X-Complex_rhown_9x9_depth_in_1_2 0.81 0.59 0.79 0.77 0.71 0.48 0.76 0.70 0.83 0.73
C2RCC_rhown_3x3_depth_in_1_2 0.81 0.60 0.81 0.81 0.76 0.62 0.79 0.58 0.78 0.76
C2X-Complex_rhow_15x15_depth_in_1_2 0.81 0.65 0.70 0.74 0.81 -0.44 0.75 0.80 0.77 0.78
C2X_rhow_5x5_depth_in_1_2 0.80 0.72 0.64 0.74 0.79 0.72 0.65 0.68 0.66 0.81
TABLE VII: Test RMSE (mg/m3) for depth 1-2 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
C2X-Complex_rhow_5x5_depth_in_1_2 1.60 2.70 1.53 2.05 1.82 2.28 2.09 2.20 1.62 1.75
C2X-Complex_rhown_5x5_depth_in_1_2 1.58 2.59 1.62 1.98 1.90 2.31 2.17 2.28 1.59 1.99
C2X-Complex_rhow_9x9_depth_in_1_2 1.61 2.85 1.70 2.10 1.92 2.91 2.50 1.97 1.75 1.75
C2X_rhow_3x3_depth_in_1_2 1.78 2.29 2.62 2.11 1.69 2.43 2.12 2.02 2.57 1.76
C2X-Complex_rhown_9x9_depth_in_1_2 1.85 2.72 1.94 2.05 2.28 3.06 2.08 2.33 1.75 2.22
C2X-Complex_rhow_3x3_depth_in_1_2 1.93 2.81 1.92 1.97 2.09 2.87 2.26 2.35 1.76 2.06
C2X-Complex_rhown_3x3_depth_in_1_2 1.81 2.72 1.76 2.11 2.18 2.45 2.18 2.31 1.98 2.05
C2RCC_rhown_3x3_depth_in_1_2 1.86 2.69 1.83 1.86 2.10 2.62 1.97 2.75 2.00 2.06
C2X-Complex_rhow_15x15_depth_in_1_2 1.84 2.52 2.34 2.15 1.84 5.11 2.12 1.89 2.05 2.01
C2X_rhow_5x5_depth_in_1_2 1.88 2.25 2.55 2.16 1.97 2.26 2.51 2.40 2.23 1.85
Tables VIII and IX present results for the depth range of 2 to 3 meters. Interestingly, at this depth the unprocessed datasets (TOA) outperformed the processed alternatives, with performance improving as the aggregation window increased (the best result corresponding to 15×15, followed by 9×9, 5×5, and 3×3). The ensemble achieved the highest $R^{2}$ (0.81) with an RMSE of 1.61 mg/m3. Notably, KNN also performed exceptionally well at this depth.
TABLE VIII: Test $R^{2}$ for depth 2-3 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
TOA_15x15_depth_in_2_3 0.73 0.40 0.81 0.78 0.65 0.21 0.63 0.62 0.52 0.66
TOA_9x9_depth_in_2_3 0.71 0.39 0.79 0.78 0.67 0.27 0.65 0.63 0.52 0.63
TOA_3x3_depth_in_2_3 0.74 0.43 0.76 0.71 0.72 0.35 0.72 0.65 0.48 0.69
TOA_5x5_depth_in_2_3 0.75 0.41 0.76 0.75 0.72 0.36 0.69 0.62 0.50 0.73
C2RCC_rhown_5x5_depth_in_2_3 0.71 0.58 0.73 0.70 0.67 0.61 0.69 0.64 0.65 0.68
C2X-Complex_rhown_9x9_depth_in_2_3 0.73 0.44 0.61 0.58 0.65 0.41 0.68 0.64 0.68 0.66
C2X-Complex_rhow_5x5_depth_in_2_3 0.72 0.16 0.71 0.59 0.63 0.24 0.64 0.55 0.72 0.65
C2X-Complex_rhown_5x5_depth_in_2_3 0.68 0.18 0.60 0.58 0.58 -0.48 0.68 0.54 0.72 0.61
C2RCC_rhow_3x3_depth_in_2_3 0.68 0.54 0.68 0.67 0.59 0.55 0.69 0.61 0.63 0.59
C2X_rhow_9x9_depth_in_2_3 0.69 0.62 0.62 0.58 0.64 0.64 0.63 0.63 0.67 0.65
TABLE IX: Test RMSE (mg/m3) for depth 2-3 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
TOA_15x15_depth_in_2_3 1.88 2.82 1.61 1.72 2.15 3.25 2.21 2.24 2.54 2.13
TOA_9x9_depth_in_2_3 1.97 2.86 1.67 1.70 2.10 3.12 2.16 2.23 2.53 2.23
TOA_5x5_depth_in_2_3 1.82 2.80 1.78 1.81 1.93 2.92 2.04 2.24 2.59 1.88
C2X-Complex_rhown_9x9_depth_in_2_3 1.81 2.61 2.16 2.25 2.06 2.67 1.97 2.09 1.96 2.03
TOA_3x3_depth_in_2_3 1.87 2.76 1.81 1.95 1.94 2.95 1.94 2.17 2.63 2.02
C2RCC_rhown_5x5_depth_in_2_3 1.86 2.27 1.82 1.91 2.00 2.17 1.95 2.08 2.06 1.97
C2X-Complex_rhow_5x5_depth_in_2_3 1.85 3.18 1.89 2.22 2.12 3.04 2.09 2.32 1.83 2.05
C2X-Complex_rhown_5x5_depth_in_2_3 1.96 3.16 2.19 2.26 2.26 4.23 1.98 2.36 1.84 2.18
C2RCC_rhow_3x3_depth_in_2_3 1.95 2.36 1.98 1.99 2.24 2.34 1.93 2.19 2.12 2.23
C2X_rhow_9x9_depth_in_2_3 1.93 2.14 2.13 2.26 2.09 2.08 2.10 2.12 2.01 2.05
At the 3-4 m depth, a considerable deterioration in performance was observed, shown in Tables X and XI. $R^{2}$ decreased by nearly 0.2 compared to the previous depth. The best-performing models differed depending on whether $R^{2}$ or RMSE was considered; however, C2X-Complex_rhow_5x5 provided consistently strong results across both metrics. Regarding the models, Random Forest outperforms in most cases, achieving a decent $R^{2}$ of 0.65 and an exceptional RMSE of 1.48 mg/m3, close to what was achieved on the surface.
@@
TABLE X: Test $R^{2}$ for depth 3-4 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
TOA_15x15_depth_in_3_4 0.66 0.50 0.58 0.62 0.59 0.38 0.61 0.61 0.47 0.57
TOA_3x3_depth_in_3_4 0.59 0.46 0.55 0.61 0.64 0.40 0.60 0.66 0.45 0.61
C2X-Complex_rhow_5x5_depth_in_3_4 0.59 0.50 0.52 0.57 0.62 0.41 0.59 0.65 0.51 0.63
TOA_5x5_depth_in_3_4 0.58 0.44 0.50 0.59 0.64 0.45 0.58 0.63 0.42 0.62
C2X-Complex_rhown_5x5_depth_in_3_4 0.60 0.56 0.57 0.57 0.59 0.34 0.55 0.63 0.51 0.61
TOA_9x9_depth_in_3_4 0.61 0.46 0.63 0.62 0.59 0.38 0.62 0.61 0.47 0.58
C2X-Complex_rhow_15x15_depth_in_3_4 0.49 0.60 0.43 0.42 0.46 0.41 0.62 0.55 0.54 0.50
C2X-Complex_rhow_9x9_depth_in_3_4 0.55 0.56 0.51 0.55 0.57 0.46 0.61 0.62 0.58 0.60
TOA_1x1_depth_in_3_4 0.54 0.38 0.44 0.60 0.49 0.18 0.54 0.55 0.44 0.46
C2X-Complex_rhown_9x9_depth_in_3_4 0.59 0.58 0.50 0.52 0.54 0.37 0.57 0.59 0.55 0.58
TABLE XI: Test RMSE (mg/m3) for depth 3-4 meters
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
C2X-Complex_rhow_5x5_depth_in_3_4 1.60 1.76 1.72 1.63 1.54 1.91 1.60 1.48 1.74 1.52
C2X-Complex_rhown_5x5_depth_in_3_4 1.58 1.67 1.64 1.64 1.60 2.02 1.68 1.52 1.75 1.56
C2X-Complex_rhow_9x9_depth_in_3_4 1.67 1.65 1.74 1.67 1.63 1.83 1.55 1.54 1.63 1.57
C2X-Complex_rhow_15x15_depth_in_3_4 1.80 1.60 1.91 1.92 1.86 1.95 1.56 1.69 1.72 1.80
C2X-Complex_rhown_9x9_depth_in_3_4 1.61 1.62 1.77 1.72 1.70 1.98 1.64 1.61 1.68 1.61
TOA_15x15_depth_in_3_4 1.74 2.10 1.93 1.83 1.91 2.34 1.86 1.86 2.16 1.94
TOA_3x3_depth_in_3_4 1.90 2.18 1.99 1.85 1.79 2.31 1.89 1.74 2.20 1.86
TOA_5x5_depth_in_3_4 1.92 2.23 2.10 1.91 1.78 2.21 1.93 1.81 2.27 1.84
TOA_9x9_depth_in_3_4 1.86 2.19 1.81 1.83 1.91 2.35 1.84 1.86 2.16 1.94
TOA_1x1_depth_in_3_4 2.01 2.35 2.22 1.87 2.13 2.70 2.02 2.00 2.22 2.19
@@
Overall, results indicate that the C2X-Complex processor consistently provides the best performance across most depths, confirming that its design for complex waters is well suited to the conditions of the Mar Menor. The only exception is observed at 2-3 m depth, where unprocessed TOA reflectances yielded better results. Furthermore, model accuracy is clearly depth-dependent: predictions are strongest near the surface, while performance systematically deteriorates with depth. At greater depths, none of the processing strategies, whether C2RCC, C2X, or C2X-Complex, achieve outstanding results, highlighting the intrinsic limitations of satellite-based reflectance for capturing chlorophyll variability in deeper layers of the water column. Additionally, larger window aggregations appear to more accurately capture the actual reflectance behavior. These aggregations have been shown to overcome single-pixel variability, particularly in 5x5 and 9x9 configurations, based on results across all depths.
@@
The models and datasets are selected based on both $R^{2}$ and RMSE, and they are:
  • Depth 0-1 meters: XGB with C2X-Complex_rhow_9x9.
  • Depth 1-2 meters: CAT with C2X-Complex_rhow_5x5.
  • Depth 2-3 meters: KNN with TOA_15x15.
  • Depth 3-4 meters: RF with C2X-Complex_rhow_5x5.
@@
The ensemble was excluded from this selection because training all the models and making inferences with them is costly, since this process must be repeated for each of the approximately 1.2 million pixels that make up the Mar Menor area. This decision was also made because simpler models yielded results quite close to those of the ensemble and to simplify the models’ use for those interested.
@@
Therefore, four models are trained and then saved into joblib files, along with a metadata file and the features required to run the model in JSON format.
@@
The final step consisted of performing inference with the trained models over all pixels of the Mar Menor. To achieve this, a GeoJSON file was created to delimit the lagoon area. The workflow then followed the same approach used to extract reflectances at buoy locations, but this time a mask derived from the GeoJSON was applied to select all water pixels. This resulted in a large CSV file containing approximately 1.2 million rows. The same preprocessing pipeline used during training was replicated to ensure that the dataset contained the required features. The models were then applied to predict Chl-a concentration at each depth, generating a second CSV file with the predictions. Finally, these predictions were converted into GeoTIFFs to visualize Chl-a maps at different depths. Examples of these maps are shown in Figures 3 and 4.

### v2

@@ 5.1 Results
When comparing the performance obtained in the upper layers with the studies summarized in Table 2 our metrics demonstrate a high degree of competitiveness with previous findings for the Mar Menor. However, it is worth noting that our dataset spans nearly a decade, covers the widest range of Chl-a concentrations reported so far, offers the highest spatial resolution, and includes a complete validation framework that takes into consideration the inherent variability of the dataset by repeting the experiments with 100 different seeds. These factors make the results particularly robust, albeit at the expense of a more complex and computationally demanding processing workflow.
Table 6: Test RMSLE for depth 1-2 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
C2X-Complex_rhow_15x15_depth_in_1_2 0.365 ± 0.048 0.385 ± 0.048 0.392 ± 0.055 0.394 ± 0.053 0.429 ± 0.049 0.375 ± 0.042 0.380 ± 0.050
C2X-Complex_rhow_9x9_depth_in_1_2 0.382 ± 0.047 0.398 ± 0.051 0.419 ± 0.059 0.393 ± 0.053 0.454 ± 0.065 0.387 ± 0.047 0.389 ± 0.048
C2X-Complex_rhown_9x9_depth_in_1_2 0.392 ± 0.047 0.403 ± 0.052 0.425 ± 0.064 0.412 ± 0.048 0.442 ± 0.064 0.395 ± 0.049 0.406 ± 0.043
C2X-Complex_rhow_5x5_depth_in_1_2 0.395 ± 0.047 0.409 ± 0.052 0.418 ± 0.054 0.426 ± 0.050 0.463 ± 0.060 0.393 ± 0.038 0.412 ± 0.048
C2X-Complex_rhow_3x3_depth_in_1_2 0.394 ± 0.042 0.412 ± 0.049 0.420 ± 0.051 0.422 ± 0.051 0.466 ± 0.050 0.409 ± 0.035 0.416 ± 0.046
C2X-Complex_rhown_3x3_depth_in_1_2 0.395 ± 0.045 0.412 ± 0.051 0.426 ± 0.051 0.419 ± 0.058 0.454 ± 0.055 0.413 ± 0.038 0.410 ± 0.052
C2RCC_rhown_3x3_depth_in_1_2 0.400 ± 0.041 0.405 ± 0.043 0.404 ± 0.046 0.425 ± 0.046 0.423 ± 0.049 0.408 ± 0.042 0.419 ± 0.044
C2X-Complex_rhown_5x5_depth_in_1_2 0.402 ± 0.045 0.413 ± 0.051 0.424 ± 0.053 0.432 ± 0.051 0.450 ± 0.047 0.401 ± 0.042 0.421 ± 0.048
C2X_rhow_5x5_depth_in_1_2 0.407 ± 0.041 0.418 ± 0.042 0.422 ± 0.040 0.427 ± 0.052 0.462 ± 0.056 0.406 ± 0.035 0.424 ± 0.046
C2X_rhow_3x3_depth_in_1_2 0.422 ± 0.041 0.427 ± 0.044 0.433 ± 0.045 0.442 ± 0.050 0.471 ± 0.052 0.421 ± 0.036 0.440 ± 0.047
@@
Table 7: Test $R^{2}$ for depth 1-2 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
C2X-Complex_rhow_15x15_depth_in_1_2 0.759 ± 0.097 0.729 ± 0.111 0.714 ± 0.142 0.670 ± 0.190 0.684 ± 0.128 0.728 ± 0.104 0.715 ± 0.123
C2X-Complex_rhow_9x9_depth_in_1_2 0.756 ± 0.104 0.723 ± 0.116 0.673 ± 0.148 0.715 ± 0.148 0.613 ± 0.202 0.740 ± 0.085 0.728 ± 0.126
C2X-Complex_rhow_5x5_depth_in_1_2 0.748 ± 0.096 0.711 ± 0.130 0.664 ± 0.173 0.669 ± 0.166 0.625 ± 0.173 0.747 ± 0.085 0.696 ± 0.126
C2RCC_rhown_3x3_depth_in_1_2 0.743 ± 0.115 0.730 ± 0.113 0.732 ± 0.129 0.666 ± 0.181 0.727 ± 0.131 0.722 ± 0.115 0.681 ± 0.158
C2X-Complex_rhow_3x3_depth_in_1_2 0.742 ± 0.096 0.690 ± 0.144 0.662 ± 0.147 0.646 ± 0.187 0.639 ± 0.173 0.706 ± 0.093 0.663 ± 0.135
C2X-Complex_rhown_5x5_depth_in_1_2 0.732 ± 0.102 0.704 ± 0.123 0.659 ± 0.158 0.661 ± 0.160 0.653 ± 0.177 0.717 ± 0.100 0.673 ± 0.129
C2X-Complex_rhown_9x9_depth_in_1_2 0.724 ± 0.117 0.707 ± 0.122 0.663 ± 0.167 0.645 ± 0.184 0.667 ± 0.198 0.728 ± 0.096 0.671 ± 0.140
C2X-Complex_rhown_3x3_depth_in_1_2 0.725 ± 0.118 0.682 ± 0.138 0.634 ± 0.170 0.655 ± 0.161 0.644 ± 0.196 0.671 ± 0.117 0.670 ± 0.136
C2X_rhow_5x5_depth_in_1_2 0.675 ± 0.111 0.661 ± 0.098 0.647 ± 0.110 0.619 ± 0.146 0.600 ± 0.122 0.648 ± 0.087 0.608 ± 0.129
C2X_rhow_3x3_depth_in_1_2 0.638 ± 0.140 0.630 ± 0.104 0.619 ± 0.102 0.600 ± 0.156 0.580 ± 0.139 0.595 ± 0.085 0.575 ± 0.141
Tables 8 and 9 present results for the depth range of 2 to 3 meters, where TOA reflectances unexpectedly outperformed atmospherically corrected datasets. This could be attributed to the fact that atmospheric correction processors, such as C2RCC, rely on pretrained neural networks that may introduce spectral artifacts or over-correct signal attenuation in turbid and shallow, waters where optical behavior of light is not linear with depth [soriano-gonzalezCombinationC2RCCProcessors2022]. At these depths, the signal-to-noise ratio is critical and the inherent non-linearities of many ML models allow them to extract biological information directly from TOA data, bypassing the potential information loss or biases introduced during the water-leaving reflectance estimation process [gomezNewApproachMonitor2021]. Furthermore, the fact that performance improved with larger aggregation windows (peaking at $\times15$) suggests that at intermediate depths, where the signal from the water column is weaker and more susceptible to noise, spatial smoothing is particularly relevant. A $\times15$ pixel window ($\times150$ m) helps to average out local surface glint and sensor noise, capturing a more stable representative reflectance of the deeper layers that individual pixels cannot resolve, and provides the model with a more robust feature set for predicting Chl-a in the transition zone of the water column. Once more, CatBoost outperformed the other models, achieving a RMSLE of 0.38 and $R^{2}$ above 0.7. Interestingly, SVR provided the best results for both metrics in most scenarios with C2X-Complex, and this behavior was also observed at the deepest layer.
Table 8: Test RMSLE for depth 2-3 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
TOA_15x15_depth_in_2_3 0.378 ± 0.035 0.398 ± 0.038 0.384 ± 0.046 0.419 ± 0.043 0.443 ± 0.047 0.432 ± 0.037 0.415 ± 0.047
TOA_9x9_depth_in_2_3 0.388 ± 0.034 0.404 ± 0.039 0.401 ± 0.043 0.426 ± 0.046 0.436 ± 0.048 0.431 ± 0.037 0.419 ± 0.049
TOA_5x5_depth_in_2_3 0.393 ± 0.035 0.408 ± 0.038 0.429 ± 0.038 0.427 ± 0.045 0.443 ± 0.058 0.440 ± 0.038 0.423 ± 0.038
TOA_3x3_depth_in_2_3 0.409 ± 0.035 0.419 ± 0.040 0.432 ± 0.041 0.432 ± 0.045 0.462 ± 0.058 0.450 ± 0.040 0.426 ± 0.041
C2RCC_rhown_5x5_depth_in_2_3 0.424 ± 0.041 0.418 ± 0.040 0.436 ± 0.044 0.436 ± 0.044 0.455 ± 0.046 0.430 ± 0.042 0.430 ± 0.042
C2X_rhow_9x9_depth_in_2_3 0.430 ± 0.041 0.433 ± 0.038 0.441 ± 0.040 0.456 ± 0.044 0.464 ± 0.045 0.422 ± 0.032 0.453 ± 0.041
C2X-Complex_rhown_9x9_depth_in_2_3 0.426 ± 0.045 0.432 ± 0.047 0.455 ± 0.057 0.441 ± 0.045 0.466 ± 0.058 0.425 ± 0.046 0.438 ± 0.045
C2X-Complex_rhow_5x5_depth_in_2_3 0.430 ± 0.038 0.441 ± 0.044 0.455 ± 0.046 0.456 ± 0.044 0.490 ± 0.056 0.431 ± 0.039 0.445 ± 0.038
C2X-Complex_rhown_5x5_depth_in_2_3 0.436 ± 0.039 0.445 ± 0.042 0.458 ± 0.047 0.460 ± 0.042 0.480 ± 0.043 0.433 ± 0.039 0.452 ± 0.038
C2RCC_rhow_3x3_depth_in_2_3 0.447 ± 0.041 0.446 ± 0.042 0.452 ± 0.047 0.471 ± 0.043 0.472 ± 0.042 0.440 ± 0.041 0.468 ± 0.0447
@@
Table 9: Test $R^{2}$ for depth 2-3 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
TOA_15x15_depth_in_2_3 0.723 ± 0.091 0.700 ± 0.062 0.716 ± 0.095 0.636 ± 0.117 0.649 ± 0.096 0.558 ± 0.068 0.608 ± 0.134
TOA_9x9_depth_in_2_3 0.706 ± 0.091 0.695 ± 0.060 0.691 ± 0.090 0.639 ± 0.110 0.637 ± 0.127 0.562 ± 0.071 0.620 ± 0.120
TOA_5x5_depth_in_2_3 0.697 ± 0.089 0.686 ± 0.056 0.653 ± 0.082 0.631 ± 0.118 0.639 ± 0.110 0.550 ± 0.066 0.610 ± 0.113
TOA_3x3_depth_in_2_3 0.692 ± 0.078 0.675 ± 0.058 0.639 ± 0.082 0.634 ± 0.111 0.607 ± 0.145 0.526 ± 0.068 0.623 ± 0.111
C2RCC_rhown_5x5_depth_in_2_3 0.663 ± 0.085 0.681 ± 0.075 0.650 ± 0.099 0.625 ± 0.119 0.650 ± 0.097 0.660 ± 0.076 0.641 ± 0.099
C2X-Complex_rhown_5x5_depth_in_2_3 0.652 ± 0.087 0.641 ± 0.085 0.600 ± 0.120 0.607 ± 0.103 0.598 ± 0.115 0.670 ± 0.073 0.596 ± 0.086
C2X-Complex_rhown_9x9_depth_in_2_3 0.635 ± 0.099 0.641 ± 0.089 0.585 ± 0.144 0.598 ± 0.124 0.601 ± 0.133 0.665 ± 0.079 0.591 ± 0.102
C2X-Complex_rhow_5x5_depth_in_2_3 0.664 ± 0.085 0.645 ± 0.085 0.599 ± 0.117 0.610 ± 0.105 0.571 ± 0.143 0.658 ± 0.079 0.605 ± 0.097
C2RCC_rhow_3x3_depth_in_2_3 0.634 ± 0.095 0.643 ± 0.078 0.633 ± 0.113 0.578 ± 0.116 0.625 ± 0.090 0.653 ± 0.081 0.562 ± 0.100
C2X_rhow_9x9_depth_in_2_3 0.644 ± 0.091 0.646 ± 0.071 0.637 ± 0.079 0.577 ± 0.106 0.628 ± 0.111 0.628 ± 0.074 0.569 ± 0.098
At the 3-4 m depth, a considerable deterioration in performance was observed, shown in Tables 10 and 11. $R^{2}$ decreased around 0.1 compared to the previous depth, and RMSLE increased to 0.4 approximately. The observed performance drop could be driven by the attenuation coefficient of the water in the Mar Menor. In turbid Case 2 waters, the downwelling irradiance and the upwelling radiance are exponentially attenuated with depth due to absorption by phytoplankton, non-algal particles, and colored dissolved organic matter (CDOM). The other factor for this may be the influence of the bottom reflectance introducing noise and interfering with the real signal. The best-performing models were CatBoost once more, the ensemble, and SVR. As in the previous depth, TOA reflectances metrics surpassed those from C2X-Complex, with the $\times15$ aggregation window being the best one.
Table 10: Test RMSLE for depth 3-4 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
TOA_15x15_depth_in_3_4 0.392 ± 0.035 0.406 ± 0.036 0.397 ± 0.041 0.437 ± 0.046 0.451 ± 0.044 0.428 ± 0.034 0.421 ± 0.044
TOA_9x9_depth_in_3_4 0.396 ± 0.033 0.407 ± 0.036 0.406 ± 0.041 0.438 ± 0.045 0.448 ± 0.046 0.426 ± 0.035 0.425 ± 0.042
TOA_5x5_depth_in_3_4 0.408 ± 0.037 0.414 ± 0.034 0.438 ± 0.036 0.447 ± 0.042 0.461 ± 0.047 0.434 ± 0.033 0.434 ± 0.040
TOA_3x3_depth_in_3_4 0.424 ± 0.035 0.427 ± 0.033 0.439 ± 0.036 0.468 ± 0.042 0.477 ± 0.056 0.436 ± 0.034 0.444 ± 0.041
C2X-Complex_rhow_15x15_depth_in_3_4 0.432 ± 0.044 0.440 ± 0.043 0.448 ± 0.050 0.457 ± 0.050 0.485 ± 0.045 0.433 ± 0.042 0.443 ± 0.044
TOA_1x1_depth_in_3_4 0.434 ± 0.036 0.440 ± 0.030 0.456 ± 0.038 0.474 ± 0.038 0.484 ± 0.053 0.448 ± 0.032 0.452 ± 0.037
C2X-Complex_rhow_9x9_depth_in_3_4 0.447 ± 0.048 0.450 ± 0.045 0.464 ± 0.052 0.461 ± 0.050 0.495 ± 0.047 0.442 ± 0.044 0.452 ± 0.044
C2X-Complex_rhown_9x9_depth_in_3_4 0.452 ± 0.050 0.453 ± 0.045 0.472 ± 0.050 0.474 ± 0.053 0.496 ± 0.054 0.448 ± 0.044 0.456 ± 0.044
C2X-Complex_rhow_5x5_depth_in_3_4 0.456 ± 0.044 0.462 ± 0.044 0.473 ± 0.050 0.481 ± 0.049 0.503 ± 0.051 0.460 ± 0.043 0.467 ± 0.043
C2X-Complex_rhown_5x5_depth_in_3_4 0.456 ± 0.046 0.464 ± 0.043 0.480 ± 0.051 0.479 ± 0.051 0.505 ± 0.054 0.460 ± 0.042 0.466 ± 0.040
Table 11: Test $R^{2}$ for depth 3-4 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
TOA_15x15_depth_in_3_4 0.598 ± 0.092 0.587 ± 0.083 0.565 ± 0.118 0.488 ± 0.160 0.510 ± 0.137 0.491 ± 0.100 0.480 ± 0.130
TOA_9x9_depth_in_3_4 0.587 ± 0.091 0.584 ± 0.086 0.556 ± 0.115 0.505 ± 0.143 0.501 ± 0.177 0.498 ± 0.097 0.485 ± 0.128
TOA_5x5_depth_in_3_4 0.566 ± 0.093 0.575 ± 0.078 0.494 ± 0.115 0.490 ± 0.132 0.503 ± 0.138 0.482 ± 0.097 0.476 ± 0.120
TOA_3x3_depth_in_3_4 0.538 ± 0.118 0.549 ± 0.090 0.492 ± 0.117 0.459 ± 0.154 0.473 ± 0.155 0.467 ± 0.105 0.446 ± 0.126
TOA_1x1_depth_in_3_4 0.510 ± 0.118 0.526 ± 0.082 0.464 ± 0.126 0.424 ± 0.157 0.437 ± 0.153 0.458 ± 0.103 0.417 ± 0.142
C2X-Complex_rhow_9x9_depth_in_3_4 0.474 ± 0.137 0.499 ± 0.109 0.466 ± 0.149 0.429 ± 0.176 0.448 ± 0.135 0.524 ± 0.085 0.439 ± 0.097
C2X-Complex_rhow_15x15_depth_in_3_4 0.507 ± 0.127 0.518 ± 0.108 0.492 ± 0.165 0.420 ± 0.196 0.469 ± 0.111 0.523 ± 0.091 0.443 ± 0.165
C2X-Complex_rhown_9x9_depth_in_3_4 0.435 ± 0.152 0.474 ± 0.108 0.427 ± 0.149 0.361 ± 0.195 0.443 ± 0.125 0.498 ± 0.090 0.414 ± 0.095
C2X-Complex_rhow_5x5_depth_in_3_4 0.477 ± 0.137 0.483 ± 0.118 0.447 ± 0.164 0.415 ± 0.180 0.444 ± 0.141 0.466 ± 0.110 0.404 ± 0.119
C2X-Complex_rhown_5x5_depth_in_3_4 0.461 ± 0.172 0.475 ± 0.121 0.432 ± 0.180 0.408 ± 0.190 0.430 ± 0.162 0.455 ± 0.100 0.406 ± 0.099
@@
Overall, results indicate that the C2X-Complex processor consistently provides the best performance at the most superficial depths, confirming that its design for complex waters is well suited to the conditions of the Mar Menor, while at 2-3 m and 3-4 m depths unprocessed TOA reflectances yielded better results. Furthermore, model accuracy is clearly depth-dependent: predictions are strongest near the surface, while performance systematically deteriorates with depth. At greater depths, none of the processing strategies, whether C2RCC, C2X, or C2X-Complex, achieve outstanding results, highlighting the intrinsic limitations of satellite-based reflectance for capturing chlorophyll variability in deeper layers of the water column. Additionally, larger window aggregations appear to more accurately capture the actual reflectance behavior. These aggregations have been shown to overcome single-pixel variability, particularly in $\times15$ and $\times9$ configurations, based on results across all depths. Clearly, CatBoost surpasses the other models in most cases. This superiority can be attributed to its specialized architecture, which is specifically tailored to handle categorical features [NEURIPS2018_14491b75], thereby extracting greater predictive value from the seasonal indicators added during the data preparation stage. As a tree-based ensemble, CatBoost is also robust to noisy and highly correlated predictors [hastieElementsStatisticalLearning2009], a common challenge when working with multispectral bands, where adjacent wavelengths often exhibit strong multicollinearity. This challenge is exacerbated by the addition of spectral indices, which are defined as band combinations. CatBoost’s unique "Ordered Boosting" scheme and use of symmetric trees effectively mitigate overfitting and prediction shift. This allows the model to maintain stability even when dealing with redundant or imperfect spectral data.
@@
A multi-criteria selection process was conducted for each depth to determine the most effective combination of processing methods, aggregation windows, and ML algorithms. Unlike traditional approaches that rely solely on $R^{2}$ or absolute error, this study established a weighted hierarchy of metrics, prioritizing RMSLE >> $R^{2}$ >> RMSE to ensure models are both globally accurate and robust across the observed chlorophyll-a concentration scale. The selection workflow involved the following stages: (i) The dataset-model pairs were initially ranked by their mean RMSLE in the test phase to prioritize sensitivity to initial phytoplankton surges while mitigating the impact of extreme eutrophic outliers. (ii) Statistical stability was evaluated by analyzing the standard deviation of both RMSLE and $R^{2}$ across the 100 seeds, penalizing configurations with high variance to ensure model transferability. (iii) A final combined score was calculated to identify optimal solutions through a weighted balance of performance and consistency. This score was composed of 50% mean RMSLE, 35% mean $R^{2}$, 10% RMSLE standard deviation, and 5% $R^{2}$ standard deviation, providing a refined trade-off between predictive power and reliability. (iv) A final check was performed on the top-ranked candidates by comparing training and test metrics to confirm high generalization capabilities and exclude any models exhibiting signs of overfitting. Following this procedure, the models and datasets selected to be used for inference are:
  • Depth 0-1 meters: CAT and C2X-Complex_rhow_15x15.
  • Depth 1-2 meters: CAT and C2X-Complex_rhow_15x15.
  • Depth 2-3 meters: CAT and TOA_15x15.
  • Depth 3-4 meters: CAT and TOA_15x15.
@@
The ensemble was excluded from the selection process because training all the models and making inferences with them is costly, since this process must be repeated for each of the approximately 1.2 million pixels that make up the Mar Menor area. This decision was also made because simpler models yielded results quite close or even better than those of the ensemble as well as to simplify the models’ use for those interested.
@@
Once selected, the four models were re-trained using the full available dataset (100%) to maximize predictive power for new observations,and then saved into joblib files, along with a metadata file and the features required to run the model in JSON format.
@@
The final step consisted of performing inference with the trained models over all pixels of the Mar Menor. To achieve this, a GeoJSON file was created to delimit the lagoon area. The workflow then followed the same approach used to extract reflectances at buoy locations, but this time a mask derived from the GeoJSON was applied to select all water pixels. This resulted in a large CSV file containing approximately 1.2 million rows. The same preprocessing pipeline used during training was replicated to ensure that the dataset contained the required features. The models were then applied to predict Chl-a concentration at each depth, generating a second CSV file with the predictions. Finally, these predictions were converted into GeoTIFFs to visualize Chl-a maps at different depths, overlayed with bathymetric contour lines to provide a clearer geographical and physical context of the lagoon’s basin. Examples of these maps are shown in Figures 3 and 4.